If 15.65 mL of aq. HCl is required to neutralise 20.00 mL of the sodium carbonate 

solution from question 4 above, what is the molarity of the aq. HCl? (Show all steps in 

your calculation as set out below.)

(a). Write the reaction equation for the titration.

(b). Calculate the number of moles of Na2CO3 (standard solution) used in the titration.

(c). Using the balanced reaction equation, find the number of moles of HCl that 

reacted.

(d). Calculate the molarity of the HCl solution.


1
Expert's answer
2021-04-25T06:15:29-0400

The reaction involved in this titration is

(a) "Na_2CO_3 (aq) + 2HCl \longrightarrow2NaCl+ H_2O+CO_2"

(b) mass of Na2CO3 = 2.5 g,molar mass of Na2CO3 =106

moles of Na2CO3 ="\frac{2.5}{106}=0.0235"

(c) from above reaction we can say that

1 mole Na2CO3 reacts with 2 mole HCl

so, 0.0235 mole HCl react with 0.047 mole HCl

(d) Molarity = "\frac{mole\times 1000}{volume (inmL)}"


Molarity="\frac{0.047\times 1000}{20(inmL)}" =2.35 M